OpenStudy (anonymous):

How would you solvethis by completing the square ? x^2 - 15x = -26 can someone please explain it to me? and this one -x^2 + x + 6 = 0

OpenStudy (anonymous):

\[x^2 - 15x = -26 \] is going to be annoying because \(15\) is even, but we can still complete the square, we just have to work with fractions

OpenStudy (anonymous):

half of \(15\) is \(\frac{15}{2}\) and \(\left(\frac{15}{2}\right)^2=\frac{225}{4}\) write \[(x-\frac{15}{2})^2=-26+\frac{225}{4}\] and finish that way

OpenStudy (anonymous):

is it 13 and 2 ?